### v1

@@ 5.2 Quantitative Results
We adopted Structural Similarity Index(SSIM) [23], Multi-Scale Structural Similarity (MS-SSIM), Fréchet Inception Distance (FID) [24] and Kernel Inspection Distance (KID) [25] scores to measure the similarity between ground truths and synthesized images. The ground truths were made by manually wrapping clothes over the models’ images using various imaging tools and the synthesized images were the output of the model used. The result are shown in Table  1.
@@ 5.3 User Study
Although SSIM, MS-SSIM, FID, and KID can be used to determine the quality of image synthesis, it cannot reflect the overall realism and visual quality as assessed by human evaluation. Thus we performed a user study with 60 volunteers. In order to evaluate realism, volunteers were provided with two sets of images: ground truth images (manually wrapped clothes on the human models) and the outputs generated by our model. They were asked to score based on how real the clothes looked on the person and how well the texture of clothing was preserved. Then, they were asked to independently rate the photo-realism of only our output images. The result shows that our result was 70% similar to the ground truth and 60% photo-realistic.
Table 1: Quantitative evaluation of CP-VTON+ and VTON-IT in terms of SSIM, MS-SSIM, FID and KID scores.
@@ 6 Discussion
The effectiveness of this approach has been demonstrated through experiments conducted on both male and female bodies, under different lighting conditions, and with occlusions such as hands and hair, as well as varying poses. The first stage of the human detection network plays a crucial role in enhancing the overall performance of the application by reducing the input image size for later stages and eliminating unnecessary input regions by cropping the human body. To create an accurate segmentation map that adheres to geometric principles, it is necessary to train a human body segmentation network using ground truth images that contain detailed wrist and neck regions. Problem with public dataset like LVIS [26], MS COCO [8] and, Pascal Person Part dataset [27] is improper ground truth, hence we tried different pre-trained models like CDCL [28] , GRAPHONY [29], U2NET [30] to generate ground truth but none of these generate precise body mask. So, we manually picked 6000 good images of both men and women from the FGC6 dataset and labeled them manually. We trained a Generative image translation model on conditional settings with a pair of images as input. In order to generate a pair of images, we utilize geometrical augmentation techniques on both the semantic mask generated by the segmentation network and the corresponding real RGB image.
@@ 7 Conclusion and Future Works
The paper presents an innovative approach called VTON-IT (Virtual Try-On using Image Translation) that enables the transfer of desired clothing onto a person’s image, while accounting for variations in body size, pose, and lighting conditions. The proposed method outperforms existing approaches, as evidenced by quantitative and qualitative results that demonstrate the generation of natural-looking synthesized images. The proposed architecture consists of three components: human detection, body part segmentation, and image translation network. The image translation network is trained in this paper to generate synthesized images of the same domain (sweatshirts). However, it can be adapted for cross-domain synthesis by incorporating control parameters as label features. Future work may extend this approach to different types of clothing such as trousers, shorts, shoes, and more.
Acknowledgements: The authors would like to thank IKebana Solutions LLC for providing them with constant support for this research project.

### v2

@@ 5.2 Quantitative Results
We adopted Structural Similarity Index(SSIM) [24], Multi-Scale Structural Similarity (MS-SSIM), Fréchet Inception Distance (FID) [25], and Kernel Inspection Distance (KID) [26] scores to measure the similarity between ground truths and synthesized images. The ground truths were made by manually wrapping clothes over the models’ images using various imaging tools and the synthesized images were the output of the model used. The results are shown in Table  1.
@@ 5.3 User Study
Although SSIM, MS-SSIM, FID, and KID can be used to determine the quality of image synthesis, it cannot reflect the overall realism and visual quality as assessed by human evaluation. Thus we performed a user study with 60 volunteers. To evaluate realism, volunteers were provided with two sets of images: ground truth images (manually wrapped clothes on the human models) and the outputs generated by our model. They were asked to score based on how real the clothes looked on the person and how well the texture of the clothing was preserved. Then, they were asked to independently rate the photo realism of only our output images. The result shows that our result was 70% similar to the ground truth and 60% photo-realistic.
Table 1: Quantitative evaluation of CP-VTON+ and VTON-IT in terms of SSIM, MS-SSIM, FID, and KID scores.
@@ 6 Discussion
The effectiveness of this approach has been demonstrated through experiments conducted on both male and female bodies, under various lighting conditions, and with occlusions such as hands and hair, as well as varying poses. The initial stage of the human detection network plays a crucial role in enhancing the overall performance of the application by reducing the input image size for subsequent stages and eliminating unnecessary input regions through cropping the human body.
To create an accurate segmentation map adhering to geometric principles, it is essential to train a human body segmentation network using ground truth images containing detailed wrist and neck regions. However, existing public datasets like LVIS [27], MS COCO [8], and Pascal Person Part dataset [28] suffer from improper ground truth annotations. Therefore, we attempted to use various pre-trained models like CDCL [29], GRAPHONY [30], and U2NET [31] to generate ground truth annotations, but none of these produced precise body masks. Consequently, we manually curated 6000 high-quality images of both men and women from the FGC6 dataset and labeled them manually.
We trained a generative image translation model in conditional settings with a pair of images as input. To generate a pair of images, we utilized geometric augmentation techniques on both the semantic mask generated by the segmentation network and the corresponding real RGB image.
@@ 7 Conclusion and Future Works
The paper introduces an innovative approach named VTON-IT (Virtual Try-On using Image Translation) that facilitates the transfer of desired clothing onto a person’s image while accommodating variations in body size, pose, and lighting conditions. This method surpasses existing approaches, as evidenced by both quantitative and qualitative results showcasing the generation of natural-looking synthesized images. The proposed architecture comprises three components: human detection, body part segmentation, and an image translation network. While this paper focuses on training the image translation network to generate synthesized images within the same domain (specifically, sweatshirts), it can be adapted for cross-domain synthesis by incorporating control parameters as label features. Future work could extend this approach to different types of clothing such as trousers, shorts, shoes, and beyond.
Acknowledgements: The authors extend their gratitude to IKebana Solutions LLC for their constant support throughout this research project.